A cart is at x = 5.25m and time t = 0. The cart accelerates at 

4.15m/s^2. If The speed of the cart at t=0 is 3m/s, find the position 

of the cart at t =2s and also determine where the cart is when it 

reaches a speed of 5m/s.


1
Expert's answer
2021-04-28T13:03:13-0400
"x(2)=5.25+3(2)+0.5(4.15)(2)^2=19.55\ m"

"5^2-3^2=2(4.15)x\\x=1.93\ m"